Huang Automotive is presently operating at 75% of capacity. The company recently received an offer from a Korean truck manufacturer to purchase 28,000 units of a power steering system component for $198 per unit. Peter Wu, vice-president of sales, notes that although there will be an additional $2.00 shipping cost for each component, he thinks that accepting the order will get the company's "foot in the door" of an expanding international market.

To determine variable and fixed costs, Huang's accountant used the high-low method with the following production and cost information for the last two years:

196,000 234,000
Direct Material costs 15,974,000 19,071,000
Direct Labor Costs 5,880,000 7,020,000
Overhead Costs 22,192,000 24,168,000
Selling & Admin Costs 7,356,000 7,774,000
Total Costs 51,402,000 58,033,000
Total Costs Per Unit 262.26 248.00

T.J. Chan, vice-president of engineering, feels that any new market should first show its profitability and that the $198 per unit offer is not only below the regular $250 selling price, but it's below the unit cost of the component. She also points out that there will be additional setup costs of $300,000 and that Huang will have to lease some special equipment for $290,000.

REQUIRED [6 tries]
1. Using the high-low method to determine cost behavior, what would the expected profit be on the special order (use a negative sign for a loss)
Answer #1

FIXED COSTS are constant and they will not change with acceptance of order

so they are not relevant for decision making

we will find out fixed and variable cost.

activity LOW a activity HIGH b change a-b variable cost [change in cost/change in activity]
activity 196000 234000 38000
Direct Material costs 15,974,000 19,071,000 3097000 81.5[3097000/38000]
Direct Labor Costs 5,880,000 7,020,000 1140000 30$[1140000/38000]
Overhead Costs 22,192,000 24,168,000 1976000 52 [1976000/38000]
Selling & Admin Costs 7,356,000 7,774,000 418000 11[418000/38000]

material cost per unit =$81.5*196000units = 15974000$ WHICH IS EQUAL TO TOTAL COST

so whole cost is variable 81.5$

labor cost = 30*196000=5880000

so whole cost is variable $30

overhead cost = 196000*52

=10192000

total cost = fixed cost+variable cost

22192000= fixed + 10192000

fixed cost =22192000-10192000

=12000000

selling and administration = 11*196000=2156000

7356000 =fixed + 2156000

fixed = 5200000$

cost for 28000unitS

SPECIAL ORDER PROFIT

Revenue 5,544,000 [198*28000]
direct material (2282000) [81.5*28000]
labor (840000) [30*28000]
overhead [variable part only] (1456000) [52*28000]
selling and administrative [variable part only] (308000) [11*28000]
Additional shipping cost (56000) [2*28000]
set up cost (300000)
special equipment (290000)
net income(Loss) 12000

profit=12000$